# Robust Learning in Bayesian Parallel Branching Graph Neural Networks: The Narrow Width Limit

### v2

@@ Abstract
The infinite width limit of random neural networks is known to result in Neural Networks as Gaussian Process (NNGP) (Lee et al. (2018)), characterized by task-independent kernels. It is widely accepted that larger network widths contribute to improved generalization (Park et al. (2019)). However, this work challenges this notion by investigating the narrow width limit of the Bayesian Parallel Branching Graph Neural Network (BPB-GNN), an architecture that resembles residual GCN. We demonstrate that when the width of a BPB-GNN is significantly smaller compared to the number of training examples, each branch exhibits more robust learning due to a symmetry breaking of branches in kernel renormalization. Surprisingly, the performance of a BPB-GNN in the narrow width limit is generally superior to or comparable to that achieved in the wide width limit in bias-limited scenarios. Furthermore, the readout norms of each branch in the narrow width limit are mostly independent of the architectural hyperparameters but generally reflective of the nature of the data. Our results characterize a newly defined narrow-width regime for parallel branching networks in general.
@@ 1 Introduction
The study of neural network architectures has seen substantial growth, particularly in understanding how network width impacts learning and generalization. In general, wider networks are believed to perform better (Allen-Zhu et al. (2019); Jacot et al. (2018); Gao et al. (2024)). However, this work challenges the prevailing assumption by exploring the narrow width limit of Bayesian Parallel Branching Graph Neural Networks (BPB-GNNs), an architecture inspired by residual GCN networks (Chen et al. (2020a; 2022)). We show theoretically and empirically that narrow-width networks can perform better than their wider counterparts due to a symmetry-breaking effect in kernel renormalization, in bias-limited scenarios. This paper presents a detailed theoretical analysis of BPB-GNNs in the narrow-width regime, highlighting realistic conditions under which these networks demonstrate robust learning and comparable generalization.
@@
    We introduce a novel yet simple GCN architecture with parallel independent branches, and derive the exact generalization error for node regression in the statistical limit as the sample size $P\to\infty$ and network width $N\to\infty$, with their ratio a finite number $\alpha=P/N$, in the over-parameterized regime.
@@ 2 Related works
Our work follows a long tradition of mathematical analysis of infinitely-wide neural networks (Neal (2012); Jacot et al. (2018); Lee et al. (2018); Bahri et al. (2024)), resulting in NTK or NNGP kernels. Recently, such analysis has been extended to structured neural networks, including GNNs (Du et al. (2019); Walker & Glocker (2019)). However, they do not provide an analysis of feature learning in which the kernel depends on the tasks.
@@
There is a long line of works that theoretically analyze the expressiveness (Xu et al. (2018); Geerts & Reutter (2022)) and generalization performance (Tang & Liu (2023); Garg et al. (2020); Aminian et al. (2024)) of GNN. However, it is challenging to calculate the dependence of generalization errors on tasks. To our knowledge, our work is first to provide a tight bound of the generalization error for GNN with residual-like structures. The architecture closest to our linear BPB-GCN is the linearly decoupled GCN proposed by Cong et al. (2021); however, the overall readout vector is shared for all branches, which will not result in kernel renormalization for different branches.
3 BPB-GNN
3.1 Parallel branching GNN architecture
We are motivated to study the parallel branching networks as they resemble residual blocks in commonly used architectures (Kipf & Welling (2016); He et al. (2016); Chen et al. (2020b)) and are easy to study analytically with our Bayesian framework. Given graph $G=(A,X)$, where $A$ is the adjacency matrix and $X$ the node feature matrix, the final readout for node $\mu$ is a scalar $f^{\mu}(G;\Theta)$ that depends on the graph and network parameters $\Theta$. The parallel branching GNN is an ensemble of GNN branches, where each branch operates independently with no weight sharing. In this work, we focus on the simple setup of branches made of linear GCN with one hidden layer, but with different number of convolutions $A$ [^1] In this paper, the convolution operation is normalized as $A=D^{-1/2}(\hat{A}+I)D^{-1/2}$, where $\hat{A}$ is the original Adjacency matrix and $D$ is the degree matrix. We also use feature standardization after convolution for each branch to normalize the input on the input node features (Figure 1(a)). In this way, parallel branching GNN is analogous to GCN with residual connections, for which the final node readout can also be thought of as an ensemble of convolution layers (Veit et al. (2016)). Concretely, the overall readout $f^{\mu}(G;\Theta)$ for node $\mu$ is a sum of $L$ branch readouts $f_{l}^{\mu}$:
@@ 3.1 Parallel branching GNN architecture
Note that when $L=2$, the network reduces exactly to a 2-layer residual GCN (Chen et al. (2020c)). Here $N_{0}$ is the input dimension, $N$ the width of the hidden layer, $W^{(l)}$ the weight of the hidden layer for branch $l$ and $a^{(l)}$ the final readout weight for branch $l$. We will consider only the linear activation function for this paper and provide a summary of notations in Appendix A.1.
@@
Figure 1: Overview of the main takeaway: BPB-GNN learns robust representations for each branch at narrow width. (a) The parallel branching GNN architecture, with 2 branches. The independent branches have non-sharing weights and produce the final output $f$ as a sum of branch-level readouts $f_{l}$. (b) Student and teacher readout norms squared for wide and narrow student BPB-GNN networks. The student network with width $N$ is trained with the teacher network’s output. Histograms correspond to the samples from Hamiltonian Monte Carlo simulations and solid lines correspond to the order parameters calculated theoretically. $\sigma_{t}=\sigma_{w}=1$. At $N=4$, the HMC samples of branch readout norms squared (orange and red histograms) for the student network $\frac{\|a_{l}\|^{2}}{N}\sigma_{w}^{2}$ concentrate at their respective theoretical values $u_{l}\sigma_{w}^{2}$ and overlap with the teacher’s readout norms squared $\frac{\|A_{l}\|^{2}}{N}\sigma_{t}^{2}$ (orange and red dashed lines) for corresponding branches. At $N=1024$ the samples for the student network (blue and green histograms) concentrate at their respective theoretical values but remain far from the teacher’s values, instead approaching the GP limit $\sigma_{w}^{4}$ (blue dashed line).
@@ 3.3 Kernel renormalization and order parameters
for each branch $l$. Therefore, branches in general become more and more different as the width of the hidden layer $N$ decreases. We call this phenomenon symmetry breaking, which is discussed in Section 4.2.
@@ 3.4 Predictor statistics and generalization
Under the theoretical framework, we obtain analytically (Appendix A.3) the mean prediction and variance for a single test node $\nu$ as
$\langle f^{\nu}(G)\rangle=\sum_{l}\frac{1}{L}u_{l}k_{l,\nu}(K+TI)^{-1}Y$ (14)
$\langle\delta f_{\nu}(G)^{2}\rangle=K_{\nu,\nu}-\sum_{l^{\prime}}\frac{1}{L}u_{l^{\prime}}k_{l^{\prime},\nu}^{T}(K+TI)^{-1}\sum_{l^{\prime\prime}}\frac{1}{L}u_{l^{\prime\prime}}k_{l^{\prime\prime},\nu}$ (15)
Here
$k_{l}^{\nu}=\frac{\sigma_{w}^{2}}{N_{0}}[A^{l}XX^{T}A^{l}]|_{(P,\nu)}$ (16)
$K^{\nu,\nu}=\frac{\sigma_{w}^{2}}{N_{0}}\sum_{l}\frac{1}{L}u_{l}[A^{l}XX^{T}A^{l}]|_{(\nu,\nu)}$ (17)
is the single matrix element for the test node. We will use this theoretical prediction to calculate the generalization performance, defined by the MSE on $t$ test nodes
@@
Note that our definition of bias and variance is a statistical average over the posterior weight distribution, which is slightly different from the usual definition from GD. However, we expect that the overall generalization error to be similar to the one resulting from one random initialization, when the test sample size is large enough.
@@ 4.1 Robust learning of branches: the equipartition conjecture
$f^{*}(G;\Theta^{*})=\sum_{l}f_{l}^{*}(G;W_{l}^{*})=\sum_{l}\frac{1}{\sqrt{N_{t}L}}\sum_{i}A_{i,l}h_{i}^{l}(G;W_{l}^{*}).$ (20)
$W_{ij}^{*}\sim\mathcal{N}(0,\sigma_{t}^{2})$ and $A_{i,l}\sim\mathcal{N}(0,\beta_{l}^{2})$, where $\beta_{l}^{2}$ is the variance assigned to the readout weight for the teacher branch $l$ and $N_{t}$ is the width of the hidden layer. Similarly, the student network is given by the same architecture, with layer width $N$ and learns from $P$ node lables from the teacher $Y_{\mu=1}^{*P}=f^{*}(G,\Theta^{*})_{\mu}$ in the Bayesian regression setup of Eq. 4 with prior variance $\sigma_{w}^{2}$. We conjecture that as $\alpha=P/N\to\infty$ the posterior distribution of the student network readout vector $a_{l}$ satisfies
$\sigma_{w}^{2}\langle\|a_{l}\|^{2}\rangle=\sigma_{t}^{2}\|A_{l}\|^{2}.$ (21)
Sketch of proof:
At narrow width, the saddle equation 12 becomes $r_{l}=\mathrm{Tr}_{l}$. Now consider $(Y^{*}Y^{*T})_{\mu,\nu}=\sum_{i,j,l_{1},l_{2}}\frac{1}{N_{t}L}A_{i,l_{1}}A_{j,l_{2}}h_{i}^{l_{1},\mu}(G)h_{j}^{l_{2},\nu}(G)$; we conjecture that this quantity concentrates at its expectation value $Y^{*}Y^{*T}\approx\mathbb{E}_{a^{*},W^{*}}(YY^{T})=\sum_{l}\beta_{l}^{2}K_{l}/L$. Given this assumption, $r_{l}$ becomes
$r_{l}=Y^{*T}K^{-1}\frac{u_{l}K_{l}}{L}K^{-1}Y^{*}=\mathrm{Tr}(K^{-1}\frac{u_{l}K_{l}}{L}Y^{*}Y^{*T})\approx\mathrm{Tr}_{l}(K^{-1}\sum_{l}\beta_{l}^{2}\frac{K_{l}}{L})$ (22)
The solution that satisfies the saddle point equations is
@@
Therefore, by Eq.13, we proved the conjecture given $N_{t}$ is also large enough. We call this equipartition conjecture, as the mean-squared readout and the variance (A.4) have to exactly balance each other, which contribute to the energy term in the Hamiltonian. This is only a conjecture, as it relies on the concentration equality assumption made in the proof. However, as we demonstrate in the experiments, the equipartition conjecture holds empirically.
Furthermore, writing $u_{l}K_{l}/L\approx Y^{*}_{l}Y^{*T}$, at narrow width, the predictor statistics for each stream for the training data from the teacher becomes
$\langle f_{l}(X)\rangle=\frac{u_{l}K_{l}}{L}(K+TI)^{-1}Y^{*}\approx Y^{*}_{l},$ (24)
ie. not only do we recover the statistics for the teacher $A_{l}$’s, we have also recovered the feature learned by each branch.
@@ 4.2 Student-Teacher experiment on robust branch learning
We demonstrate this robust learning phenomenon and provide a first evidence of the equipartition conjecture with the student-teacher experiment setup introduced in the previous section. We use the contextual stochastic block model (CSBM) (Deshpande et al. (2018)) to generate the graph, where the adjacency matrix is given by a stochastic block model with two blocks, and the node feature is generated with latent vectors corresponding to the two blocks (Appendix B.1). Both student and teacher network has $L=2$ branches. We calculate the order parameters $u_{l}$’s for the student network with the saddle point equations, as well as perform Hamiltonian Monte Carlo (HMC) sampling (Appendix B.3) from the posterior distribution of the student’s network weights to generate a distribution of student readout norms squared $a_{l}^{2}$. We compare the values of the student branch readout norms to their corresponding teacher branches for the student network with either narrow or wide width. As shown in Figure 1(b), an extremely narrow student network ($N=4$) learns the teacher’s branch readout norms very robustly, while a much wider network ($N=1024$) fails to learn the teachers’ norms and approaches the GP ($N\to\infty$) limit where the two branches are indistinguishable.
@@
5 BPB-GNN on Cora
We also perform experiments on the Cora benchmark dataset (McCallum et al. (2000)) by training the BPB-GNN with binary node regression, for a range of $L,N,\sigma_{w}$ values (Appendix B.2 for details). We observe a similar narrow-to-wide width transition for the bias term. As shown in Figure 5, the bias increases with network width, transitioning to the GP regime, and we observe the trend extending to a potential narrow width limit. [^2] In this case, the narrow width limit is hard to demonstrate as the transition window is below realistic minimum of network width. Additionally, it is demonstrated that using more branches that involve higher-order convolutions improves performance.

### v3

# When narrower is better: the narrow width limit of Bayesian parallel branching neural networks

@@ Abstract
The infinite width limit of random neural networks is known to result in Neural Networks as Gaussian Process (NNGP) (Lee et al. (2017)), characterized by task-independent kernels. It is widely accepted that larger network widths contribute to improved generalization (Park et al. (2019)). However, this work challenges this notion by investigating the narrow width limit of the Bayesian Parallel Branching Neural Network (BPB-NN), an architecture that resembles neural networks with residual blocks. We demonstrate that when the width of a BPB-NN is significantly smaller compared to the number of training examples, each branch exhibits more robust learning due to a symmetry breaking of branches in kernel renormalization. Surprisingly, the performance of a BPB-NN in the narrow width limit is generally superior to or comparable to that achieved in the wide width limit in bias-limited scenarios. Furthermore, the readout norms of each branch in the narrow width limit are mostly independent of the architectural hyperparameters but generally reflective of the nature of the data. We demonstrate such phenomenon primarily for branching graph neural networks, where each branch represents a different order of convolutions of the graph; we also extend the results to other more general architectures such as the residual-MLP and show that the narrow width effect is a general feature of the branching networks. Our results characterize a newly defined narrow-width regime for parallel branching networks in general.
@@ 1 Introduction
The study of neural network architectures has seen substantial growth, particularly in understanding how network width impacts learning and generalization. In general, wider networks are believed to perform better (Allen-Zhu et al. (2019); Jacot et al. (2018); Gao et al. (2024)). However, this work challenges the prevailing assumption by exploring the narrow width limit of Bayesian Parallel Branching Neural Network (BPB-NN), an architecture inspired by neural networks with residual blocks (He et al. (2016); Chen et al. (2020a; 2022)). We show theoretically and empirically that narrow-width BPB-NNs can perform better than their wider counterparts due to the symmetry-breaking effect in kernel renormalization, in bias-limited scenarios. We present a detailed analysis of BPB-NNs primarily by exploring the Bayesian Parallel Branching Graph Convolution Networks (BPB-GCN), while extending the results to other more general architectures such as the residual-MLP in the appendix.
@@
    We introduce a novel yet simple graph neural architecture with parallel independent branches, and derive the exact generalization error for node regression in the statistical limit as the sample size $P\to\infty$ and network width $N\to\infty$, with their ratio a finite number $\alpha=P/N$, in the over-parameterized regime.
@@
  4. 4.
    We further show that this narrow-width effect is a general feature of Bayesian parallel branching neural networks (BPB-NNs), with the residual-MLP architecture as an example.
@@ 2 Related works
Our work follows a long tradition of mathematical analysis of infinitely-wide neural networks (Neal (2012); Jacot et al. (2018); Lee et al. (2017); Bahri et al. (2024)), resulting in NTK or NNGP kernels. Recently, such analysis has been extended to structured neural networks, including GCNs (Du et al. (2019); Walker & Glocker (2019); Huang et al. (2021)). However, they do not provide an analysis of feature learning in which the kernel depends on the tasks.
@@
There is a long line of works that theoretically analyze the expressiveness (Xu et al. (2018); Geerts & Reutter (2022)) and generalization performance (Tang & Liu (2023); Garg et al. (2020); Aminian et al. (2024)) of GCN. However, it is challenging to calculate the dependence of generalization errors on tasks. In particular, the PAC-Bayes approach Liao et al. (2020); Ju et al. (2023) results in generalization bounds that are too large and that can be only computed with norms of learned weights. To our knowledge, our work is first to decompose the generalization error into bias and variance a priori (not dependent on learned weights) for linear GCNs with residual-like structures. The architecture closest to our linear BPB-GCN is the linearly decoupled GCN proposed by Cong et al. (2021); however, the overall readout vector is shared for all branches, which will not result in kernel renormalization for different branches.
3 BPB-GCN
3.1 Parallel branching GCN architecture
We are motivated to study the parallel branching networks as they resemble residual blocks in commonly used architectures (Kipf & Welling (2016); He et al. (2016); Chen et al. (2020b)) and are easy to study analytically with our Bayesian framework. We primarily focus on the graph setting (BPB-GCN) in the main sections, and discuss the more general case in appendix B.
Given graph $G=(A,X)$, where $A$ is the adjacency matrix and $X$ the node feature matrix, the final readout for node $\mu$ is a scalar $f^{\mu}(G;\Theta)$ that depends on the graph and network parameters $\Theta$. The parallel branching GCN is an ensemble of GCN branches, where each branch operates independently with no weight sharing. In this work, we focus on the simple setup of branches made of linear GCN with one hidden layer, but with different number of convolutions $A$ [^1] In this paper, the convolution operation is normalized as $A=D^{-1/2}(\hat{A}+I)D^{-1/2}$, where $\hat{A}$ is the original Adjacency matrix and $D$ is the degree matrix. We also use feature standardization after convolution for each branch to normalize the input on the input node features (Figure 1(a)). In this way, parallel branching GCN is analogous to GCN with residual connections, for which the final node readout can also be thought of as an ensemble of convolution layers (Veit et al. (2016)). Concretely, the overall readout $f^{\mu}(G;\Theta)$ for node $\mu$ is a sum of $L$ branch readouts $f_{l}^{\mu}$:
@@ 3.1 Parallel branching GCN architecture
Note that when $L=2$, the network reduces exactly to a 2-layer residual GCN (Chen et al. (2020c)). Here $N_{0}$ is the input dimension, $N$ the width of the hidden layer, $W^{(l)}$ the weight of the hidden layer for branch $l$, $A^{l}$ the $l$th power of the adjacency matrix and $a^{(l)}$ the final readout weight for branch $l$. We will consider only the linear activation function for this paper and provide a summary of notations in Appendix A.1.
@@
Figure 1: Overview of the main takeaway: BPB-GCN learns robust representations for each branch at narrow width. (a) The parallel branching GCN architecture, with 2 branches. The independent branches have non-sharing weights and produce the final output $f$ as a sum of branch-level readouts $f_{l}$. (b) Student and teacher readout norms squared for wide and narrow student BPB-GCN networks. The student network with width $N$ is trained with the teacher network’s output. Histograms correspond to the samples from Hamiltonian Monte Carlo simulations and solid lines correspond to the order parameters calculated theoretically. $\sigma_{t}=\sigma_{w}=1$. At $N=4$, the HMC samples of branch readout norms squared (orange and red histograms) for the student network $\frac{\|a_{l}\|^{2}}{N}\sigma_{w}^{2}$ concentrate at their respective theoretical values $u_{l}\sigma_{w}^{2}$ and overlap with the teacher’s readout norms squared $\frac{\|A_{l}\|^{2}}{N}\sigma_{t}^{2}$ (orange and red dashed lines) for corresponding branches. At $N=1024$ the samples for the student network (blue and green histograms) concentrate at their respective theoretical values but remain far from the teacher’s values, instead approaching the GP limit $\sigma_{w}^{4}$ (blue dashed line).
@@ 3.3 Kernel renormalization and order parameters
for each branch $l$, where we use the bracket notation to represent the expectation value over the posterior distribution. Therefore, branches in general become more and more different as the width of the hidden layer $N$ decreases. We call this phenomenon symmetry breaking, which is discussed in Section 4.2.
@@ 3.4 Predictor statistics and generalization
Under the theoretical framework, we obtain analytically (Appendix A.3) the mean predictor $\langle f^{\nu}(G)\rangle$ and variance $\langle\delta f_{\nu}(G)^{2}\rangle$ for a single test node $\nu$ as Eq.54 and Eq.55, respectively. We use this to calculate the generalization performance, defined by the MSE on $t$ test nodes
@@
Note that our definition of bias and variance is a statistical average over the posterior weight distribution, which is slightly different from the usual definition from GD.
@@ 4.1 Robust learning of branches: the equipartition conjecture
$f^{*}(G;\Theta^{*})=\sum_{l}f_{l}^{*}(G;W_{l}^{*})=\sum_{l}\frac{1}{\sqrt{N_{t}L}}\sum_{i}a^{*}_{i,l}h_{i}^{l}(G;W_{l}^{*}).$ (16)
$W_{ij,l}^{*}\sim\mathcal{N}(0,\sigma_{t}^{2})$ and $a^{*}_{i,l}\sim\mathcal{N}(0,\beta_{l}^{2})$, where $\beta_{l}^{2}$ is the variance assigned to the readout weight for the teacher branch $l$ and $N_{t}$ is the width of the hidden layer. Similarly, the student network is given by the same architecture, with layer width $N$ and learns from $P$ node labels from the teacher $Y_{\mu}=f^{*}_{\mu}(G,\Theta^{*})$ in the Bayesian regression setup of Eq. 4 with prior variance $\sigma_{w}^{2}$. We state that at $T=0$ temperature, in the limit $N_{t}\to\infty$ and $\alpha=P/N\to\infty$, there exists a solution to the saddle point equations 12
$u_{l}=\sigma_{t}^{2}\beta_{l}^{2}/\sigma_{w}^{2}.$ (17)
Proof:
At the narrow width limit $\alpha\to\infty$, the saddle equation 12 becomes $r_{l}=\mathrm{Tr}_{l}$. Now we calculate
$Y^{*}_{\mu}Y^{*T}_{\mu}=\sum_{i,j=1}^{N_{t}}\sum_{l_{1},l_{2}}\frac{1}{N_{t}L}a^{*}_{i,l_{1}}a^{*}_{j,l_{2}}h_{i}^{l_{1},\mu}(G;W_{l})h_{j}^{l_{2},\nu}(G;W_{l})$ (18)
by law of large numbers, this quantity concentrates at its expectation value
$Y^{*}Y^{*T}=\lim_{N_{t}\to\infty}\mathbb{E}_{a^{*},W^{*}}(YY^{T})=\sum_{l}\frac{\beta_{l}^{2}\sigma_{t}^{2}}{N_{0}L}(A^{l}XX^{T}A^{l})|_{P}=\frac{\beta_{l}^{2}}{L}K_{l}(\sigma_{t}^{2}),$ (19)
where $K_{l}(\sigma^{2})$represents the GP input kernel with prior variance $\sigma^{2}$. Therefore, $r_{l}$ becomes
$r_{l}=Y^{*T}K^{-1}\frac{u_{l}K_{l}}{L}K^{-1}Y^{*}=\mathrm{Tr}(K^{-1}\frac{u_{l}K_{l}}{L}Y^{*}Y^{*T})=\mathrm{Tr}_{l}((\sum_{l}u_{l}K_{l}(\sigma_{w}^{2})/L)^{-1}\sum_{l}\beta_{l}^{2}\frac{K_{l}}{L})$ (20)
Thus there exists a solution that satisfies the saddle point euations
@@
Furthermore, by Eq.13, we show that the student branch norms learn exactly the teacher branch norms, ie. $\langle a_{l}^{2}\rangle\sigma_{w}^{2}=a^{*2}_{l}\beta_{l}^{2}$ We call this equipartition theorem, as the mean-squared readout and the variance (A.4) have to exactly balance each other, which contribute to the energy term in the Hamiltonian. Although the theorem is proven for the BPB-GCN architecture, we show that the result is general in Appendix B.1.
@@ 4.2 Student-Teacher experiment on robust branch learning
We demonstrate this robust learning phenomenon and provide a first evidence of the equipartition conjecture with the student-teacher experiment setup introduced in the previous section. We use the contextual stochastic block model (CSBM) (Deshpande et al. (2018)) to generate the graph, where the adjacency matrix is given by a stochastic block model with two blocks, and the node feature is generated with latent vectors corresponding to the two blocks (Appendix C.1). Both student and teacher network has $L=2$ branches. We calculate the order parameters $u_{l}$’s for the student network with the saddle point equations, as well as perform Hamiltonian Monte Carlo (HMC) sampling (Appendix C.5) from the posterior distribution of the student’s network weights to generate a distribution of student readout norms squared $a_{l}^{2}$. We compare the values of the student branch readout norms to their corresponding teacher branches for the student network with either narrow or wide width. As shown in Figure 1(b), an extremely narrow student network ($N=4$) learns the teacher’s branch readout norms very robustly, while a much wider network ($N=1024$) fails to learn the teachers’ norms and approaches the GP ($N\to\infty$) limit where the two branches are indistinguishable.
@@
5 BPB-GCN on Cora
We also perform experiments on the Cora benchmark dataset (McCallum et al. (2000)) by training the BPB-GCN with binary node regression, for a range of $L,N,\sigma_{w}$ values (Appendix C.2 for details). We observe a similar narrow-to-wide width transition for the bias term. As shown in Figure 5, the bias increases with network width, transitioning to the GP regime, and we observe the trend extending to a potential narrow width limit. [^2] In this case, the narrow width limit is hard to demonstrate as the transition window is below realistic minimum of network width. Additionally, it is demonstrated that using more branches that involve higher-order convolutions improves performance.